# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

### v1

@@ Abstract.
The ability to research and synthesize knowledge is central to human expertise and progress. An emerging class of systems promises these exciting capabilities through generative research synthesis, performing retrieval over the live web and synthesizing many discovered sources into long-form, cited summaries. However, evaluating such systems remains an open challenge: existing question-answering benchmarks focus on short-form factual responses, while expert-curated datasets risk staleness and data contamination. Both fail to capture the complexity and evolving nature of real research synthesis tasks. In this work, we introduce DeepScholar-bench, a live benchmark and holistic, automated evaluation framework designed to evaluate generative research synthesis. DeepScholar-bench draws queries from recent, high-quality ArXiv papers and focuses on a real research synthesis task: generating the related work sections of a paper by retrieving, synthesizing, and citing prior research. We develop an automated evaluation framework that holistically assesses performance across three key dimensions, knowledge synthesis, retrieval quality, and verifiability, using metrics that show strong agreement with expert human judgments. We also develop DeepScholar-base, a reference pipeline for generative research synthesis, implemented efficiently using the LOTUS API. Using the DeepScholar-bench framework, we perform a systematic evaluation of prior open-source systems, search AI’s with open-source and strong proprietary models, OpenAI’s DeepResearch, and DeepScholar-base. We find that DeepScholar-base establishes a strong baseline, attaining competitive or higher performance than prior open-source systems, search AI’s and OpenAI’s DeepResearch. We also find that DeepScholar-bench remains far from saturated, with no system exceeding a score of across all metrics. These results underscore both the difficulty and the importance of DeepScholar-bench as a foundation for progress toward AI systems capable of generative research synthesis. We make our benchmark code and data available at https://github.com/guestrin-lab/deepscholar-bench.
1. Introduction
@@
A crucial underpinning of human knowledge and innovation is the ability of human experts to research and synthesize known facts and new findings to support others in comprehending, verifying and building upon existing work. Recently, new systems designed for generative research synthesis offer exciting capabilities and promises to automate challenging research synthesis tasks, which often require laborious hours of literature search, reading, and writing from human experts. These systems include a wide array of both commercial DeepResearch offerings, including ones from OpenAI (noa, [n.d.]g), Gemini (noa, [n.d.]c), Anthropic (noa, [n.d.]b), Grok (noa, [n.d.]e), and Perplexity (noa, [n.d.]h), as well as open-source methods, such as STORM (Shao et al., 2024), DeepResearcher (Zheng et al., 2025), and OpenScholar (Asai et al., 2024). These systems push the frontier of AI capabilities, demonstrating promising performance on existing factuality and question-answering benchmarks (Wei et al., 2024; Krishna et al., 2025; Mialon et al., 2023; Wei et al., 2025).
@@
Yet, as this new class of systems emerges, a key question remains: how should we benchmark and evaluate generative research synthesis? The progress of these systems requires benchmarks that carefully evaluate their critical capabilities. Specifically, these systems provide three core functions: (1) retrieval, typically from a large, complex, and constantly-evolving corpus, such as the live web, to collect key information (2) knowledge synthesis, to generate coherent long-form answers that surface key facts, integrating general knowledge and findings from many retrieved sources, and (3) verifiability, providing citations that allow readers to trace each stated claim in the synthesized answer to a reputable source from the retrieved set. The ideal benchmark must holistically evaluate across all three of these dimensions, while providing a realistic and challenging research synthesis task, which the community can reliably and repeatedly benchmark on over time.
@@
Unfortunately, existing benchmarks fall-short of these goals. Many prior works evaluate generative research synthesis systems using existing question answering benchmarks (Wei et al., 2025, 2024; Mialon et al., 2023; Krishna et al., 2025; Wu et al., 2025; Wadden et al., 2020; Jin et al., 2019; Yang et al., 2018; Joshi et al., 2017; Kwiatkowski et al., 2019; Ho et al., 2020; Trivedi et al., 2022; Lee et al., 2023), which do not reflect realistic research synthesis tasks and instead focus on questions with short-form, easily-verifiable answers, making them severely limited for this setting. These question-answering benchmarks do not capture the complexity of long-form answers synthesized from many sources, a key component of research synthesis. To address this limitation, several recent works (Asai et al., 2024; Zheng et al., 2024; noa, [n.d.]k) instead leverage expert-curated datasets with open-ended research questions and exemplar answers. Unfortunately, these benchmarks quickly become stale and outdated as new information emerges. Furthermore, these datasets risk data contamination as new models are trained on snapshots of the web, including public datasets. The prohibitive expense of curating, maintaining, and updating expert-curated benchmarks further limits their utility towards realistic, scalable evaluation.
To address this gap, we introduce DeepScholar-Bench, a live benchmark dataset and holistic evaluation framework tailored to tracking progress on generative research synthesis, as well as DeepScholar-Base, a reference method, which we hope will further future research. Our benchmark provides a live dataset with real research synthesis tasks, as well as a holistic, automated evaluation strategy. We design our benchmark dataset based on a fundamental research synthesis task reflected in thousands of research papers each year, where peer-reviewed manuscripts must concisely summarize existing work within a subfield, detailing and citing key prior works, before discussing new contributions. Specifically, the DeepScholar-bench task requires a system to generate a related work section for a given academic paper by retrieving sources from the Web and summarizing key prior work within a field of study. This task allows us to assess the key capabilities of research synthesis while ensuring that our evaluation reflects real, high-quality research. To source realistic and challenging queries, we develop an automated data pipeline that curates high-quality, recent ArXiv papers across diverse scientific domains. We plan to continuously re-run our data pipeline to provide an evolving dataset of live queries.
To assess performance on DeepScholar-bench, we develop an automated evaluation framework, which holistically measures the performance of system answers across the three key dimensions: knowledge synthesis, retrieval quality and verifiability. Evaluating the accuracy of our long-form synthesis task is inherently difficult since each query admits many possible answers, lacking a straightforward notion of correctness. Moreover, developing an automated evaluation requires reliable metrics that exhibit high agreement with expert human annotators, another significant challenge. To address these challenges, our holistic evaluation assesses each system response across seven key metrics (Table 1), which permit many possible correct answers, often leveraging human-written exemplars from our dataset. Specifically, on the knowledge synthesis dimension, we evaluate a generated answer’s Organization, using pairwise comparisons to human exemplars, and Nugget Coverage, assessing the efficiency of the generated response in capturing key information and facts. To assess retrieval quality, we measure the Relevance Rate of retrieved sources, the Document Importance of sources, according to citation counts of each reference, and Reference Coverage, by assessing the generated report’s coverage of notable important references recovered from the human-written exemplar. To assess the Verifiability of each report, we measure its Citation Precision, whether each citation supports the given claim, and Citation Coverage, measuring whether each claim is fully supported by the cited sources. Our human agreement study validates that these automated metrics are effective, demonstrating strong agreement between our model-based judges and expert human annotators.
@@
| Organization | assesses organization and coherence of system answer |
@@
Using the DeepScholar-bench framework, we systematically evaluate the performance of existing systems, including open-source research synthesis systems, search AI’s with strong proprietary models, OpenAI DeepResearch, and DeepScholar-base. Figure 1 demonstrates that all of these existing methods exhibit significant opportunity for improvement, with no baseline exceeding a score of across all metrics. Furthermore, on several key metrics, including Nugget Coverage, Reference Coverage and Document Importance, each evaluated method’s performance remains well below , reflecting the inherent difficulty of the DeepScholar-bench task, which requires systems to navigate the live web, reasoning about the relevance and importance of documents as well as surfacing key facts into a concise final answer. Notably, OpenAI’s DeepResearch offers strong performance relative to other baselines, outperforming many prior methods on knowledge synthesis and retrieval quality, with scores of on Nugget Coverage, on Reference Coverage and on Document Importance; however, it struggles to provide strong verifiability relative to many other methods. We also find that DeepScholar-base, a relatively simple reference pipeline implemented efficiently using the LOTUS API (noa, 2025b; Patel et al., 2025), consistently improves upon performance of prior open-source systems and search AI’s, as shown in Figure 1, as well as achieving competitive performance and up to higher verifiability compared to OpenAI’s DeepResearch, as shown in Figure 1. Nevertheless, DeepScholar-bench remains far from saturated, representing exciting opportunities for further work. We hope that our benchmark framework and reference pipeline support the progress of new systems, and we believe that resolving DeepScholar-bench represents a critical milestone towards more capable AI systems.
Overall, our main contributions are the following:
We propose DeepScholar-bench, a live benchmark dataset with real research synthesis tasks and an automated evaluation that holistically assesses performance across three key dimensions: knowledge synthesis, retrieval quality, and verifiability. Our analysis demonstrates that our DeepScholar-bench evaluation metrics are robust, exhibiting strong human agreement scores.
We develop DeepScholar-base, a reference pipeline for generative research synthesis, providing a strong baseline with competitive or higher performance compared to prior open-source systems, search AI’s, and OpenAI’s DeepResearch.
We perform a systematic evaluation of existing baselines on DeepScholar-bench, including open-source systems, search AI’s, OpenAI’s DeepResearch, and DeepScholar-base. We find that DeepScholar-bench is far from saturated, demonstrating significant opportunity for improvement, with no baseline achieving a score greater than across all metrics.
2. The DeepScholar Dataset
We study the task of generating a related works section of an academic paper, a fundamental research synthesis task. We choose this task for two key reasons. First, this task is a real research task performed by academic experts, allowing our benchmark to reflect realistic, difficult and useful queries. Second, the online availability of diverse, high-quality academic papers allows us to develop an automated dataset construction pipeline that we can continuously run to obtain new queries over time.
We construct our dataset by scraping papers from ArXiv (noa, [n.d.]a), which continuously posts thousands of new pre-print papers across a wide array of scientific domains each week. We formalize our dataset task as follows: given a description, of a paper, the goal is to retrieve a set of relevant sources, , and generate a related works sections, , for the paper by synthesizing and citing the retrieved documents. We briefly overview our automated data collection framework (Section 2.1) and describe the dataset instantiation (Section 2.2) used in our evaluation (Section 5).
2.1. Automated Data Collection Framework
@@
Focus on recent research papers, both to provide realistic, timely benchmark queries and to control data contamination when benchmarking models trained on snapshots of the web.
@@
Control for quality of the scraped ArXiv papers and extracted data
Figure 2 provides an overview of our dataset pipeline, which collects and extracts metadata about each paper (e.g., the title, abstract, and ArXiv link), the paper’s related works section, and information on each citation from the paper’s related works section. Our data collection pipeline extracts this information by scraping and selectively filtering ArXiv papers according to a number of configured settings.
@@
Specifically, the pipeline loads papers from a list of configured ArXiv domain categories (e.g., cs.ML) and filters paper according to the configured publication-date range. To avoid possible data contamination arising from multiple ArXiv versions, some of which may have been released prior to the configured publication-date range, we exclusively include v1 ArXiv papers. To control for paper quality, our pipeline optionally provides a configuration setting which filter’s out papers which are not listed as ”accepted” or ”published” at a conference within the paper’s comment metadata, which often lists updates to the paper’s status. We also disclude papers that do not have an explicit ”Related Works” section and .bib file, containing well-formatted bibliography entries. For each paper, we then extract the Related Works section, from both the LaTex files, and PDF file, if available. We clean the extracted LaTex section to remove figures, sub-figures, labels and comments. We also extract all citations found in the related work section from the LaTex .bib file. For each citation we use the ArXiv API and the OpenAlex API to recover more detailed information, such as abstracts, authors, and links for ArXiv and non-ArXiv references respectively.
2.2. Dataset Description and Statistics
@@
We now briefly summarize the dataset we use in our evaluation in Section 5 which represents an instantiation of our automated data collection pipeline. Our datasets take ArXiv papers with a publication date between April and June, following the April 5th, 2025 release date of the Llama-4 models (noa, 2025a), the main open-source model we benchmark in our evaluation. Our dataset consist of papers scraped from a diverse set of 18 disticnt ArXiv domains, including, cs.AI, cs.CV, cs.DB, cs.LG, cs.AR, cs.CG, cs.DC, cs.DS, cs.IR. To control for quality, we filter out papers not accepted at a conference, and we additionally exclude papers with related works sections longer than 1,000 words. Our final dataset instantiation includes 63 ArXiv papers, each providing a single query and expert-written exemplar for our benchmark. We make our scripts available to allow others to configure different datasets, and we plan to update our dataset to provide a continual evaluation with recent queries. Our experiments leverage the abstract of each paper as the paper’s description , provided to each baseline system as context within the query. We analyze the human-written exemplars from our dataset, and we find that, on average, each related work section contains unique references, and we find over of all cited references on ArXiv.
3. The DeepScholar Evaluation Framework
Evaluating research synthesis systems is challenging due to the complexity of the task and the variability of possible correct answers. Research synthesis systems generate complex, long-form reports, which are difficult to evaluate and lack a notion of ”ground truth” correctness. The task we consider departs significantly from traditional question answering and RAG-based evaluations (Joshi et al., 2017; Lee et al., 2023; Jin et al., 2019; Kwiatkowski et al., 2019; Trivedi et al., 2022; Yang et al., 2018; Ho et al., 2020), requiring a carefully designed and holistic evaluation framework. An exemplar research report must retrieve important relevant sources, synthesize an informative and well-organized answer, and provide appropriate references allowing readers to verify and re-trace facts. Our holistic evaluation framework thus analyzes three key dimensions, providing an automated, scalable approach for each: knowledge synthesis (Section 3.1), retrieval quality (Section 3.2), and verifiability (Section 3.3). While our experiments in Section 5 evaluate one specific task, our evaluation framework may extend to a wide range of research synthesis tasks (Shao et al., 2024; noa, [n.d.]k; Zheng et al., 2024; Asai et al., 2024), which exhibit similar properties and challenges. We provide an overview of our evaluation framework in this section and further details and analysis in Appendix Section 8.4.
3.1. Knowledge Synthesis
We evaluate both the information content surfaced in each synthesized report and the overall organization and coherence of the report.
@@ 3.1. Knowledge Synthesis
Organization and Coherency. Our automatic evaluation adopts an LLM-as-a-judge approach to assess the organization and coherence of each system answer. We use preference-based pairwise comparisons, where the LLM-judge is presented with the details of the criteria to judge, the human-written exemplar, and a generated report and is asked to mark which is better. To avoid position bias (Li et al., 2025), we evaluate each pair twice, permuting their positions. This model-based evaluation provides scalability while also serving as a strong surrogate for human preferences (Rahmani et al., 2024b; Li et al., 2024, 2025; Arabzadeh and Clarke, 2025), which we validate in our experiments in Section 5. We report the win-rate of each system using the prompt shown in Box 8.4.3 in the Appendix.
@@
Nugget Coverage. To assess the quality of the information content presented by a generated report, we use a nugget-based evaluation. An information nugget (Pradeep et al., 2025; Upadhyay et al., 2024b; Faggioli et al., 2023; Rahmani et al., 2024a; Upadhyay et al., 2024a) is an essential fact or components relevant for an answer. The process of nuggetization involves decomposing information-dense text into essential components, which aid in evaluation. For our task, we generate nuggets from the human-written exemplar related-work section for each query, following the automated, LLM-based methodology of Pradeep et al. (Pradeep et al., 2025). In section 5 we validate that the model-based approach has strong agreement with expert annotations when labeling nuggets. For, each report, we compute the nugget coverage score, the fraction of nuggets that are present in each system answer.
3.2. Retrieval Quality

### v2

@@ Abstract
The ability to research and synthesize knowledge is central to human expertise and progress. A new class of AI systems—designed for generative research synthesis—aims to automate this process by retrieving information from the live web and producing long-form, cited reports. Yet, evaluating such systems remains an open challenge: existing question-answering benchmarks focus on short, factual answers, while expert-curated datasets risk staleness and data contamination. Neither captures the complexity and evolving nature of real research synthesis tasks. We introduce DeepScholar-bench, a live benchmark and automated evaluation framework for generative research synthesis. DeepScholar-bench draws queries and human-written exemplars from recent, high-quality ArXiv papers and evaluates a real synthesis task: generating a related work section by retrieving, synthesizing, and citing prior work. Our automated framework holistically measures performance across three key dimensions—knowledge synthesis, retrieval quality, and verifiability. To further future work, we also contribute DeepScholar-ref, a simple, open-source reference pipeline, which is implemented on the LOTUS framework and provides a strong baseline. Using DeepScholar-bench, we systematically evaluate prior open-source systems, search agents with strong models, OpenAI’s DeepResearch, and DeepScholar-ref. We find DeepScholar-bench is far from saturated: no system surpasses a geometric mean of across all metrics. These results highlight both the difficulty and importance of DeepScholar-bench as a foundation for advancing AI systems capable of generative research synthesis. We make our benchmark code and data available at https://github.com/guestrin-lab/deepscholar-bench.
1 Introduction
@@
A core foundation of human knowledge and innovation is the ability of human experts to research and synthesize known facts and new findings, enabling others to comprehend, verify and build upon prior work. Recently, systems for generative research synthesis have emerged, promising to automate tasks that produce long-form outputs (e.g., multi-page reports), which traditionally demand hours of literature searching, reading and writing by human experts. These offerings include commercial ones—from OpenAI [33], Gemini [13], Anthropic [2], Grok [59], and Perplexity [38]—as well as open-source methods, such as STORM [43], DeepResearcher [65], and OpenScholar [6]. Existing systems demonstrate promising performance on factuality and question-answering benchmarks [55, 20, 31, 56], pushing the frontier of AI capabilities.
@@
Yet, as this new class of systems emerges, a key question remains: how should we benchmark and evaluate generative research synthesis? The progress of these systems requires benchmarks that carefully evaluate their critical capabilities—specifically, three core functions: (1) retrieval, typically from a large, complex, and constantly-evolving corpus, such as the live web, to collect key information (2) knowledge synthesis, to generate coherent, long-form answers that surface key facts, integrating general knowledge and findings from many retrieved sources, and (3) verifiability, providing citations that allow readers to trace each stated claim in the synthesized answer to a reputable source from the retrieved set. The ideal benchmark must holistically evaluate across all three of these dimensions, while providing a realistic and challenging research synthesis task.
@@
Unfortunately, existing benchmarks fall short of these goals. Many prior works evaluate generative research synthesis systems using existing question answering benchmarks, which do not reflect realistic research synthesis tasks and instead focus on questions with short-form, easily-verifiable answers, making them severely limited for this setting [56, 55, 31, 20, 58, 53, 18, 61, 19, 21, 15, 49, 23]. These question-answering benchmarks do not capture the complexity of long-form answers synthesized from many sources, a key component of research synthesis. To address this limitation, several recent works instead leverage expert-curated datasets with open-ended research questions and exemplar answers [6, 66, 62, 60, 8, 45, 16]. Unfortunately, these benchmarks quickly become stale and outdated as new information emerges. Furthermore, these datasets risk data contamination as new models are trained on snapshots of the web, including public datasets. The prohibitive expense of curating, maintaining, and updating expert-curated benchmarks further limits their utility towards realistic, scalable evaluation.
In this work, we introduce DeepScholar-bench, a live benchmark and holistic, automated evaluation framework designed to evaluate generative research synthesis. DeepScholar-bench draws queries from recent, high-quality ArXiv papers and focuses on a real research synthesis task: generating the related work sections of a paper by retrieving, synthesizing, and citing prior research. We plan to provide a live benchmark, releasing updated research queries every month, and practitioners can also run our automated date pipeline to create their own dataset instantiations. Further, we develop an automated evaluation framework that leverages human-written related works extracted from each ArXiv paper and holistically assesses performance across three key dimensions— knowledge synthesis, retrieval quality, and verifiability—using metrics that show strong agreement with human judgments. To promote future work, we also develop DeepScholar-ref, a simple open-source reference pipeline for generative research synthesis implemented on the LOTUS framework [28, 37].
Using the DeepScholar-bench framework, we systematically evaluate the performance of existing systems, including open-source research synthesis systems, search agents with strong proprietary models, OpenAI DeepResearch, and DeepScholar-ref. We find that all of these existing methods exhibit significant opportunity for improvement, with no system surpassing a geometric mean of across all metrics. Furthermore, on several key metrics, including Nugget Coverage, Reference Coverage and Document Importance, each evaluated method’s performance remains well below , reflecting the inherent difficulty of the DeepScholar-bench task, which requires systems to navigate the live web, reasoning about the relevance and importance of documents as well as surfacing key facts into a cohesive final answer. Notably, OpenAI’s DeepResearch offers strong performance relative to other baselines, outperforming many prior methods on knowledge synthesis and retrieval quality, with scores of on Nugget Coverage, on Reference Coverage and on Document Importance; however, it struggles to provide strong verifiability relative to many other methods. We also find that DeepScholar-ref reference pipeline represents a strong open-source baseline offering competitive performance on most metrics and up to higher verifiability compared to OpenAI’s DeepResearch. Nevertheless, DeepScholar-bench remains far from saturated, representing exciting opportunities for further work. We hope that our benchmark framework and reference pipeline support the progress of new systems, and we believe that resolving DeepScholar-bench represents a critical milestone towards more capable AI systems.
Overall, our main contributions are the following:
We propose DeepScholar-bench, a live benchmark dataset with real research synthesis tasks and an automated, holistic evaluation.
We develop DeepScholar-ref, a simple open-source reference pipeline for generative research synthesis that attains competitive performance with open-source systems, search agents, and OpenAI’s DeepResearch across many metrics using the same models.
We perform a systematic evaluation of existing baselines on DeepScholar-bench, finding significant opportunities for improvement, with no system surpassing a geometric mean of across all metrics.
@@
| Organization & Coherency | assesses organization and coherence of system answer |
@@
2 The DeepScholar Dataset
We study the task of generating a related works section of an academic paper, a fundamental research synthesis task, for which we leverage human-written exemplars from our automated data pipeline to ground our evaluation. We source our dataset queries by scraping ArXiv papers [5] accepted at academic conferences, and we formalize our dataset task as follows: given a paper’s description , the goal is to retrieve a set of relevant sources and generate a related works sections by synthesizing and citing the retrieved documents. We briefly overview our automated data collection framework (Section 2.1) and describe the dataset instantiation (Section 2.2) used in our evaluation (Section 5). Appendix A.1 provides additional details.
2.1 Automated Data Collection Framework
@@
Focus on recent research papers, both to provide realistic, timely benchmark queries and to prevent data contamination when benchmarking models trained on snapshots of the web.
3.
@@
Control for quality of the scraped ArXiv papers and extracted data, focusing on peer-reviewed manuscripts that are accepted at academic conferences.
@@
Our data pipeline scrapes papers, filters them, and extracts content to construct datasets that include metadata about each ArXiv paper (e.g., the title, abstract, and link), the papers’ related work section, and a citation list of references found in the papers’ related work sections. Our scraper begins by loading papers from a configured set of ArXiv domains (e.g., cs.ML) and configured publication-date range. To avoid possible data contamination arising from multiple ArXiv versions, we keep only v1 ArXiv papers. To control for paper quality, our pipeline then optionally filters papers to keep only those listed as “accepted” or “published” at a conference based on the ArXiv metadata. We also exclude papers that do not have an explicit “Related Work” section and .bib file, containing well-formatted bibliography entries. Our pipeline then processes each paper to extract the Related Works section from both the LaTex files and PDF file, if available. We clean the extracted LaTex section to remove any labels and comments. We also extract all citations found in the related work section and we use the ArXiv and OpenAlex APIs [35] to recover more detailed information, such as abstracts, authors, and links for both ArXiv and non-ArXiv references.
2.2 Dataset Description and Statistics
@@
The dataset instantiation, DeepScholar-June-2025 used in our evaluation (Section 5) configures a publication date range between April and June 2025, following the April 5th, 2025 release date of Llama-4 models [27], the main open-source model we benchmark. This instantiation scrapes papers from a diverse set of 18 distinct ArXiv domains—including, cs.IR, cs.CV, cs.AI, cs.CL, cs.LG, cs.DC, cs.DB, cs.AR, cs.SD, cs.CR, cs.ET, cs.GR, cs.PL, cs.SY, cs.OS, cs.PF, cs.SE, cs.MM—and selects papers marked as accepted at a conference. We additionally exclude papers with related works sections longer than 1,000 words to control for cost. Our final dataset includes 63 ArXiv papers, each providing a single query and an extracted expert-written exemplar for our evaluation. We make our scripts available to allow others to configure different datasets, and we plan to release monthly datasets of recent queries. Our experiments leverage the abstract of each paper as the paper’s description , provided to each baseline system as context within the query. We analyze the human-written exemplars from our dataset, and we find that, on average, each related work section contains unique references, and we find over of all cited references on ArXiv. We provide additional experiments in Appendix A.3.2 on a more recent, expanded dataset DeepScholar-Nov-2025, which contains 200 queries from over 75 distinct arXiv subject areas, spanning Computer Science, Physics, Quantitative Biology, Economics, and Quantitative Finance. Our results on this dataset confirm the generalization of our benchmark and main experimental conclusions.
3 The DeepScholar Evaluation Framework
Evaluating research synthesis is inherently challenging due to the task’s complexity, and the lack of a simple, "ground truth" notion of correctness, permitting many possible answers. To address these challenges, we propose a holistic, automated evaluation that assesses 7 fine-grained metrics across three key dimensions that reflect the core capabilities of research synthesis: knowledge synthesis (Section 3.1), retrieval quality (Section 3.2), and verifiability (Section 3.3). We overview our evaluation framework in this section, and provide further details and analysis of our metrics in Appendix A.4, as well as manual validation of our LLM-based metrics in Appendix A.3.
3.1 Knowledge Synthesis
Knowledge synthesis reflects the ability of a system to generate an effective final report which surfaces key facts and information into a coherent writeup. We evaluate both the overall Organization and Coherency of each system report, as well as its factual content according to its Nugget Coverage.
@@ 3.1 Knowledge Synthesis
Organization and Coherency. We use an LLM-as-a-judge to assess organization and coherence and perform pairwise comparisons between the system generated report and the human-written exemplar from the dataset. We permute each evaluated pair of reports to avoid position bias [24], and we report the win-rate of each baseline. This model-based evaluation provides scalability while also serving as a strong surrogate for human preferences [42, 25, 24, 4], which we validate in our experiments in Appendix A.3.
@@
Nugget Coverage. To assess the quality of the information content of a generated report, we use a nugget-based evaluation. An information nugget is an essential fact or component relevant for an answer [39, 51, 50, 9, 41]. For our task, we generate nuggets from the human-written exemplar related-work section for each query, and for each generated report, we compute the nugget coverage score, the fraction of nuggets present in the answer, following the automated, LLM-based methodology of [39].
3.2 Retrieval Quality